### v2

# Lele Liu,  Yi-Zheng Fan,  Yi Wang,  Wenyan Wang

Abstract
The spread of a real symmetric matrix is defined as the difference between its largest and smallest eigenvalue. The study of graph-related matrices has attracted considerable attention, leading to a substantial body of findings. In this paper, we investigate a general spread problem related to $A_{\alpha}$-matrix of graphs. The $A_{\alpha}$-matrix of a graph $G$, introduced by Nikiforov in 2017, is a convex combinations of its diagonal degree matrix $D(G)$ and adjacency matrix $A(G)$, defined as $A_{\alpha}(G)=\alpha D(G)+(1-\alpha)A(G)$. Let $\lambda_{1}^{(\alpha)}(G)$ and $\lambda_{n}^{(\alpha)}(G)$ denote the largest and smallest eigenvalues of $A_{\alpha}(G)$, respectively. We determined the unique graph that maximizes $\lambda^{(\alpha)}_{1}(G)-\beta\cdot\lambda^{(\gamma)}_{n}(G)$ among all connected $n$-vertex graphs for sufficiently large $n$, where $0\leq\alpha<1$, $1/2\leq\gamma<1$ and $0<\beta\gamma\leq 1$. As an application, we confirm a conjecture proposed by Lin, Miao, and Guo [Linear Algebra Appl. 606 (2020) 1–22]. In addition, one of main results in [SIAM J. Discrete Math. 38 (2024) 590–608] is a simple corollary of our result by choosing $\alpha=\gamma=1/2$ and $\beta=1$.
Keywords: $A_{\alpha}$-matrix, $A_{\alpha}$-eigenvalue, spread
AMS Classification: 05C35; 05C50; 15A18
1 Introduction
Consider an $n\times n$ complex matrix $M$. The spread of $M$ is defined as the maximum distance between any two of its eigenvalues. This concept was first introduced by Mirsky [15] in 1956, with further significant results appearing in [16]. Over the years, the spread has garnered considerable interest from researchers; see, for instance, [4, 10, 18, 20].
The spread of a matrix has also drawn interest in specific cases, particularly in the context of graphs. In 2001, Gregory, Hershkowitz, and Kirkland [7] initiated the study of the spread of graphs and established several bounds on this quantity. They also conjectured that among all $n$-vertex graphs, the graph that maximizes the spread of its adjacency matrix is $K_{\lfloor 2n/3\rfloor}\vee\lceil n/3\rceil K_{1}$, which is the join of the clique on $\lfloor 2n/3\rfloor$ vertices and an independent set on $\lceil n/3\rceil$ vertices. Subsequently, many scholars contributed to this conjecture and related problems of such extremal flavor for a fixed family of graphs, such as trees [6], unicyclic graphs [22], bicyclic graphs [21]. We would like to mention that Breen, Riasanovsky, Tait, and Urschel [2] confirmed this conjecture by using techniques from the theory of graph limits and numerical analysis. For the spread of the signless Laplacian matrix of graphs, commonly referred to as the $Q$-spread, Liu et al. [14] and Oliveira et al. [19] presented some upper and lower bounds. The unique graph with maximum $Q$-spread was determined in [14] and [19], which is a union of a complete graph and an isolated vertex. Among connected graphs, the graph with the minimum $Q$-spread was independently determined by Das [3] and Fan and Fallat [5] using different proof techniques. Recently, Liu [13] determined the unique $n$-vertex connected graph attaining the maximum $Q$-spread for sufficiently large $n$. This graph, known as the kite graph $Ki_{n,n-1}$, is obtained by attaching a pendant edge to one vertex of a clique on n1n-1 vertices.
Let us begin with some definitions. Let $G$ be a simple undirected graph of order $n$. Nikiforov [17] introduced the $A_{\alpha}$-matrix of $G$, defined as a convex combinations of $D(G)$ and $A(G)$:
$A_{\alpha}(G):=\alpha D(G)+(1-\alpha)A(G),$
where $D(G)$ is the diagonal matrix whose entries are the degrees of the vertices of $G$, $A(G)$ is the adjacency matrix of $G$, and $\alpha\in[0,1]$ is a real number. Obviously, $A_{0}(G)=A(G)$, and $2A_{1/2}(G)$ is exactly the signless Laplacian matrix $Q(G)$. In this sense, researches on the family $A_{\alpha}(G)$ can underpin a unified theory of both $A(G)$ and $Q(G)$. Let $\lambda_{1}^{(\alpha)}(G)$ and $\lambda_{n}^{(\alpha)}(G)$ denote the largest and smallest eigenvalues of $A_{\alpha}(G)$, respectively. Specially, $\lambda_{1}^{(0)}(G)=\lambda_{1}(G)$ is the largest eigenvalue of $A(G)$. Moreover, $2\lambda_{1}^{(1/2)}(G)=:q_{1}(G)$ and $2\lambda_{n}^{(1/2)}(G)=:q_{n}(G)$ are the largest and smallest eigenvalues of $Q(G)$, respectively.
Generalizing the notion of spread for $A(G)$ and $Q(G)$, Lin, Miao, and Guo [11] studied the $A_{\alpha}$-spread $S_{\alpha}(G)$ of a graph $G$, which is defined as
$S_{\alpha}(G):=\lambda_{1}^{(\alpha)}(G)-\lambda_{n}^{(\alpha)}(G).$
In [11] and [12], the authors presented some lower and upper bounds on $A_{\alpha}$-spread of graphs. In particular, the following conjecture was proposed.
Conjecture 1.1 ([12]).
Let $G$ be a connected graph with $n\geq 5$ vertices. If $1/2\leq\alpha<1$, then
$S_{\alpha}(G)\leq S_{\alpha}(Ki_{n,n-1})$
with equality if and only if $G\cong Ki_{n,n-1}$.
In this paper we confirm Conjecture 1.1 for sufficiently large $n$ in a stronger form.
Theorem 1.1.
Let $0\leq\alpha<1$, $1/2\leq\gamma<1$ and $0<\beta\gamma\leq 1$. Among all connected graphs on $n$ vertices, $\lambda^{(\alpha)}_{1}(G)-\beta\cdot\lambda^{(\gamma)}_{n}(G)$ is maximized only by $Ki_{n,n-1}$ for sufficiently large $n$.
Theorem 1.1 implies that Conjecture 1.1 holds for sufficiently large $n$ by letting $\beta=1$ and $\gamma=\alpha$. Additionally, Theorem 1.1 also contributes partially to a problem posed by Breen, Riasanovsky, Tait, and Urschel [2], who asked for the maximization of the function $\beta\lambda_{1}^{(0)}(G)-(1-\beta)\lambda_{n}^{(0)}(G)$ over a specified class of graphs, where $\beta\in[0,1]$. On the other hand, by choosing $\alpha=\gamma=1/2$ and $\beta=1$, we immediately derive the following result.
Corollary 1.1 ([13]).
Among all connected graphs on $n$ vertices, $q_{1}(G)-q_{n}(G)$ is maximized only by $Ki_{n,n-1}$ for sufficiently large $n$.
Finally, we can establish a new finding by setting $\alpha=0$, $\beta=2$ and $\gamma=1/2$ in Theorem 1.1.
Corollary 1.2.
Among all connected graphs on $n$ vertices, $\lambda_{1}(G)-q_{n}(G)$ is maximized only by $Ki_{n,n-1}$ for sufficiently large $n$.
2 Preliminaries
In this section we review some facts on eigenvalue – eigenvector equation, quadratic form of matrices and necessary conclusions that will be used throughout the paper.
2.1 Terminology and notation
Let $G$ be a graph. For subset $X\subset V(G)$, the subgraph of $G$ induced by $X$ is denoted by $G[X]$, and the graph obtained from $G$ by deleting $X$ is denoted by $G\setminus X$. For subsets $X,Y\subset V(G)$, we write $E(X,Y)$ for the set of edges with one vertex in $X$ and the other in $Y$, and write $E(X)$ for the set of edges induced by $X$. For a vertex $v$ of $G$, we write $d_{G}(v)$ and $N_{G}(v)$ for the degree of $v$ and the set of neighbors of $v$ in $G$, respectively. If the underlying graph $G$ is clear from the context, simply write $d(v)$ and $N(v)$. We denote the maximum degree and minimum degree of $G$ by $\Delta(G)$ and $\delta(G)$, respectively. For graph notation and terminology undefined here, we refer the reader to [1].
2.2 Eigenvalue-eigenvector equation and quadratic form
Let $G$ be a graph of order $n$, and let $\bm{x}\in\mathbb{R}^{n}$ be an eigenvector of $A_{\alpha}(G)$ corresponding to an eigenvalue $\lambda$. The eigenvalue – eigenvector equation for $\lambda$ with respect to $u\in V(G)$ is
$\lambda x_{u}=\alpha d(u)x_{u}+(1-\alpha)\sum_{v\in N(u)}x_{v}.$ (2.1)
Clearly, the quadratic form $\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle$ can be represented as
$\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle=\sum_{uv\in E(G)}\big{(}\alpha(x_{u}-x_{v})^{2}+2x_{u}x_{v}).$ (2.2)
Additionally, the quadratic form $\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle$ can be equivalently written as
$\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle=(2\alpha-1)\sum_{u\in V(G)}d(u)x_{u}^{2}+(1-\alpha)\sum_{uv\in E(G)}(x_{u}+x_{v})^{2}.$ (2.3)
2.3 Basic properties of $A_{\alpha}(G)$
Since $A_{\alpha}(G)$ is a real symmetric matrix, Rayleigh’s principle implies the following assertion.
Proposition 2.1 ([17]).
Let $\alpha\in[0,1]$ and $G$ be a graph of order $n$. Then
$\lambda_{1}^{(\alpha)}(G)=\max_{\|\bm{x}\|_{2}=1}\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle,\quad\lambda_{n}^{(\alpha)}(G)=\min_{\|\bm{x}\|_{2}=1}\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle.$
Moreover, if $\bm{x}$ is a unit vector, then $\lambda_{1}^{(\alpha)}(G)=\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle$ if and only if $\bm{x}$ is an eigenvector to $\lambda_{1}^{(\alpha)}(G)$ and $\lambda_{n}^{(\alpha)}(G)=\langle A_{\alpha}(G)\bm{x},\bm{x}\rangle$ if and only if $\bm{x}$ is an eigenvector to $\lambda_{n}^{(\alpha)}(G)$.
Proposition 2.2 ([17]).
Let $0\leq\alpha<1$. The following conclusions hold.
  1. $(1)$
    If $\alpha\geq 1/2$, then $A_{\alpha}(G)$ is positive semidefinite.
  2. $(2)$
    $\lambda_{n}^{(\alpha)}(G)<\alpha\cdot\delta(G)$.
For $0\leq\alpha\leq 1$, $\beta>0$ and $0\leq\gamma\leq 1$, as stated in Proposition 2.1, the expression $\lambda_{1}^{(\alpha)}(G)-\beta\cdot\lambda_{n}^{(\gamma)}(G)$ can be written as
$\lambda_{1}^{(\alpha)}(G){-}\beta\cdot\lambda_{n}^{(\gamma)}(G)=\max_{\|\bm{x}\|_{2}=\|\bm{z}\|_{2}=1}\sum_{uv\in E(G)}\!\!\big{(}\alpha(x_{u}{-}x_{v})^{2}{-}\beta\gamma(z_{u}{-}z_{v})^{2}{+}2(x_{u}x_{v}{-}\beta z_{u}z_{v})\big{)}.$ (2.4)
The subsequent statement extends the inequality established by Hong, Shu, and Fang [8].
Lemma 2.1 ([9]).
Suppose that $G$ is an $n$-vertex graph with $m$ edges. Let $\delta$ and $\Delta$ be the minimum and maximum degree of $G$. If $\alpha\in[0,1)$, then
$\lambda_{1}^{(\alpha)}(G)\leq\frac{(1{-}\alpha)(\delta{-}1){+}\sqrt{(1{-}\alpha)^{2}(\delta{-}1)^{2}+4(\alpha\Delta^{2}{+}(1{-}\alpha)(2m{-}(n{-}1)\delta))}}{2}.$
3 Proof of Theorem 1.1
Throughout this section, we always assume that $\alpha\in[0,1)$, $\gamma\in[1/2,1)$, $\beta\gamma\in(0,1]$, and that $G$ is a graph that maximizes $\lambda_{1}^{(\alpha)}(G)-\beta\cdot\lambda_{n}^{(\gamma)}(G)$ among all connected $n$-vertex graphs. Let $\bm{x}$ be a nonnegative unit eigenvectors of $A_{\alpha}(G)$ corresponding to $\lambda_{1}^{(\alpha)}(G)$, and let $\bm{z}$ be a unit eigenvectors of $A_{\gamma}(G)$ corresponding to $\lambda_{n}^{(\gamma)}(G)$. We also set $\lambda_{1}^{(\alpha)}:=\lambda_{1}^{(\alpha)}(G)$ and $\lambda_{n}^{(\gamma)}:=\lambda_{n}^{(\gamma)}(G)$ for short.
Let us begin with a few basic properties that will be used repeatedly.
3.1 Preliminary lemmas
Lemma 3.1.
$\lambda_{1}^{(\alpha)}>n-2-\beta\gamma$ and $0\leq\lambda_{n}^{(\gamma)}<2/\beta$.
Proof.
For the kite graph $Ki_{n,n-1}$, according to Proposition 2.2(2), $\lambda_{n}^{(\gamma)}(Ki_{n,n-1})<\gamma$. In light of the maximality of $\lambda_{1}^{(\alpha)}-\beta\cdot\lambda_{n}^{(\gamma)}$, we conclude that
$\lambda_{1}^{(\alpha)}-\beta\cdot\lambda_{n}^{(\gamma)}$ $\geq\lambda_{1}^{(\alpha)}(Ki_{n,n-1})-\beta\cdot\lambda_{n}^{(\gamma)}(Ki_{n,n-1})$
$>\lambda_{1}^{(\alpha)}(K_{n-1})-\beta\gamma$
$=n-2-\beta\gamma.$
Since $\lambda_{n}^{(\gamma)}\geq 0$ for $\gamma\in[1/2,1)$, we have $\lambda_{1}^{(\alpha)}\geq\lambda_{1}^{(\alpha)}-\beta\cdot\lambda_{n}^{(\gamma)}>n-2-\beta\gamma$.
On the other hand, it follows from $\lambda_{1}^{(\alpha)}\leq\Delta(G)\leq n-1$ that
$n-2-\beta\gamma<\lambda_{1}^{(\alpha)}-\beta\cdot\lambda_{n}^{(\gamma)}\leq(n-1)-\beta\cdot\lambda_{n}^{(\gamma)},$
which yields that $\lambda_{n}^{(\gamma)}<(1+\beta\gamma)/\beta\leq 2/\beta$. This completes the proof of the lemma. ∎
Lemma 3.1 implies that $\lambda_{1}^{(\alpha)}>n-2-\beta$. By employing Lemma 2.1, we arrive at the subsequent assertion.
Lemma 3.2.
$2e(G)>n^{2}-\frac{3-\alpha+2\beta}{1-\alpha}n$.
Proof.
By Lemma 2.1 and $\Delta(G)\leq n-1$ we see $\lambda_{1}^{(\alpha)}\leq f(\delta(G))/2$, where $f(x)$ is given by
$f(x):=(1{-}\alpha)(x{-}1){+}\sqrt{(1{-}\alpha)^{2}(x{-}1)^{2}{+}4\big{[}\alpha(n{-}1)^{2}{+}(1{-}\alpha)(2e(G){-}(n{-}1)x)\big{]}}.$
It is claimed that $f(x)$ is non-increasing in $x$. Indeed, we have
$\frac{f^{\prime}(x)}{1-\alpha}$ $=1-\frac{2(n-1)-(1-\alpha)(x-1)}{\sqrt{(1{-}\alpha)^{2}(x{-}1)^{2}{+}4\big{[}\alpha(n{-}1)^{2}{+}(1{-}\alpha)(2e(G){-}(n{-}1)x)\big{]}}}$
$=\frac{f(x)-2(n-1)}{\sqrt{(1{-}\alpha)^{2}(x{-}1)^{2}{+}4\big{[}\alpha(n{-}1)^{2}{+}(1{-}\alpha)(2e(G){-}(n{-}1)x)\big{]}}}$
$\leq 0,$
the last inequality is due to $e(G)\leq\binom{n}{2}$. It follows that
$\lambda_{1}^{(\alpha)}\leq\frac{f(1)}{2}=\sqrt{\alpha(n-1)^{2}+(1-\alpha)(2e(G)-n+1)}.$
On the other hand, as shown in Lemma 3.1, $\lambda_{1}^{(\alpha)}>n-2-\beta$. Therefore,
$\sqrt{\alpha(n-1)^{2}+(1-\alpha)(2e(G)-n+1)}>n-2-\beta.$
By rearranging, we obtain
$2(1-\alpha)\cdot e(G)$ $>(1-\alpha)n^{2}-(3-\alpha+2\beta)n+\beta^{2}+4\beta+3$
$>(1-\alpha)n^{2}-(3-\alpha+2\beta)n.$
This completes the proof. ∎
Lemma 3.3.
For any vertices $u$ and $v$, if $\alpha(x_{u}-x_{v})^{2}-\beta\gamma(z_{u}-z_{v})^{2}+2(x_{u}x_{v}-\beta z_{u}z_{v})>0$, then $u$ and $v$ are adjacent; if $\alpha(x_{u}-x_{v})^{2}-\beta\gamma(z_{u}-z_{v})^{2}+2(x_{u}x_{v}-\beta z_{u}z_{v})<0$ and GuvG-uv is connected, then $u$ and $v$ are non-adjacent.
Proof.
According to (2.4), if $\alpha(x_{u}-x_{v})^{2}-\beta\gamma(z_{u}-z_{v})^{2}+2(x_{u}x_{v}-\beta z_{u}z_{v})>0$ and $u$, $v$ are non-adjacent, then
$~\big{(}\lambda_{1}^{(\alpha)}(G+uv)-\beta\cdot\lambda_{n}^{(\gamma)}(G+uv)\big{)}-\big{(}\lambda_{1}^{(\alpha)}(G)-\beta\cdot\lambda_{n}^{(\gamma)}(G)\big{)}$
$\geq$ $~\alpha(x_{u}-x_{v})^{2}-\beta\gamma(z_{u}-z_{v})^{2}+2(x_{u}x_{v}-\beta z_{u}z_{v})>0,$
a contradiction which implies that $u$ and $v$ are adjacent. Likewise, we can show that if $\alpha(x_{u}-x_{v})^{2}-\beta\gamma(z_{u}-z_{v})^{2}+2(x_{u}x_{v}-\beta z_{u}z_{v})<0$ and GuvG-uv is connected, then $u$ and $v$ are non-adjacent. ∎
3.2 Understanding the structure
Fix a sufficiently small constant $\varepsilon>0$, whose value will be chosen later. Let
$S:=\Big{\{}v\in V(G):|z_{v}|<\frac{\varepsilon}{\sqrt{n}}\Big{\}},\hskip 8.53581ptT:=\Big{\{}v\in V(G):x_{v}<\frac{1}{2\sqrt{n}}\Big{\}},$
and $L:=V(G)\setminus S$. We further partition $L$ into two subsets:
$B:=\Big{\{}v\in L:z_{v}>0\Big{\}},\hskip 8.53581ptC:=\Big{\{}v\in L:z_{v}<0\Big{\}}.$
We first prove that $L$ has bounded size.
Lemma 3.4.
$|L|=o(n)$.
Proof.
We assume by contradiction that $|L|=\Omega(n)$. To finish the proof we shall prove $\lambda_{n}^{(\gamma)}=\Omega(n)$, and therefore get a contradiction to $\lambda_{n}^{(\gamma)}<(1+\beta\gamma)/\beta$.
To this end, note that $L=B\cup C$. Hence, at least one of $B$ and $C$ has size $\Omega(n)$. If $|B|=\Omega(n)$, we have $|E(B)|=\Omega(n^{2})$. Otherwise, if $|E(B)|=o(n^{2})$, then
$e(G)<\binom{n}{2}-\bigg{(}\binom{|B|}{2}-o(n^{2})\bigg{)}.$
This is a contradiction to Lemma 3.2. Likewise, if $|C|=\Omega(n)$, we also have $|E(C)|=\Omega(n^{2})$. Hence, $|E(B)\cup E(C)|=\Omega(n^{2})$.
Since $\bm{z}$ is a unit eigenvector corresponding to $\lambda_{n}^{(\gamma)}$, by (2.3) we see
$\lambda_{n}^{(\gamma)}$ $=(2\gamma-1)\sum_{u\in V(G)}d(u)z_{u}^{2}+(1-\gamma)\sum_{uv\in E(G)}(z_{u}+z_{v})^{2}$
$\geq(1-\gamma)\sum_{uv\in E(G)}(z_{u}+z_{v})^{2}$
$=(1-\gamma)\sum_{uv\in E(G)}(z_{u}^{2}+z_{v}^{2})+2(1-\gamma)\sum_{uv\in E(G)}z_{u}z_{v}$
$\geq 2(1-\gamma)\sum_{uv\in E(G)}(|z_{u}z_{v}|+z_{u}z_{v})$
$=2(1-\gamma)\sum_{uv\in E(G),\,z_{u}z_{v}>0}(|z_{u}z_{v}|+z_{u}z_{v})$
$\geq 2(1-\gamma)\sum_{uv\in E(B)\cup E(C)}(|z_{u}z_{v}|+z_{u}z_{v}).$
Recall that $|z_{v}|\geq\frac{\varepsilon}{\sqrt{n}}$ for each $v\in B\cup C$, we have $\lambda_{n}^{(\gamma)}=\Omega(n)$, a contradiction completing the proof. ∎
Applying Lemma 3.4, we will demonstrate that $G$ necessarily contains a vertex of small degree.
Lemma 3.5.
The graph $G$ has at least one vertex with degree $o(n)$.
Proof.
Assume by contradiction that $d(v)=\Omega(n)$ for each $v\in V(G)$. Let $v_{0}$ be a vertex such that $|z_{v_{0}}|=\max\{|z_{w}|:w\in V(G)\}$. Without loss of generality, we can assume $z_{v_{0}}\leq-n^{-1/2}$. Consider the following eigenvalue – eigenvector equation with respect to $v_{0}$,
$0\leq\big{(}\lambda_{n}^{(\gamma)}-\gamma d(v_{0})\big{)}z_{v_{0}}=(1-\gamma)\sum_{u\in N(v_{0})}z_{u}.$
Since $\lambda_{n}^{(\gamma)}=O(1)$, $d(v_{0})=\Omega(n)$ and $z_{v_{0}}\leq-n^{-1/2}$, we have
$\sum_{u\in N(v_{0})}z_{u}=\frac{\lambda_{n}^{(\gamma)}-\gamma d(v_{0})}{1-\gamma}\cdot z_{v_{0}}=\Omega(\sqrt{n}).$
On the other hand, by Cauchy–Schwarz inequality and the fact $\|\bm{z}\|_{2}=1$, we find that
$\sum_{u\in N(v_{0})}z_{u}\leq\sum_{u\in N(v_{0})\,\cap L}|z_{u}|+\sum_{u\in N(v_{0})\setminus L}|z_{u}|\leq\sqrt{|L|}+\varepsilon\sqrt{n},$
which yields that $|L|=\Omega(n)$ for sufficiently small $\varepsilon$. This is a contradiction to Lemma 3.4. ∎
Next, we will show that $G$ has exactly one vertex of small degree. Before continuing, we need the following results.
Lemma 3.6.
For each $v\in V(G)$, $x_{v}<\frac{1}{\sqrt{n}}+O(n^{-3/2})$.
Proof.
By the eigenvalue – eigenvector equation and Cauchy–Schwarz inequality we have
$\big{(}\lambda_{1}^{(\alpha)}-\alpha d(v)\big{)}x_{v}=(1-\alpha)\sum_{u\in N(v)}x_{u}\leq(1-\alpha)\sqrt{d(v)},$
which, together with Lemma 3.1 and $d(v)\leq n-1$, gives
$x_{v}\leq\frac{(1-\alpha)\sqrt{d(v)}}{\lambda_{1}^{(\alpha)}-\alpha d(v)}<\frac{(1-\alpha)\sqrt{n}}{(1-\alpha)n-4}=\frac{1}{\sqrt{n}}+O\Big{(}\frac{1}{n^{3/2}}\Big{)},$
as desired. ∎
Lemma 3.7.
$|T|=O(1)$.
Proof.
By Lemma 3.6 and $\|\bm{x}\|_{2}=1$, we deduce that
$1$ $=\sum_{u\in T}x_{u}^{2}+\sum_{u\in V(G)\setminus T}x_{u}^{2}$
$<\frac{|T|}{4n}+(n-|T|)\bigg{(}\frac{1}{\sqrt{n}}+O\Big{(}\frac{1}{n^{3/2}}\Big{)}\bigg{)}^{2}$
$=1-\frac{3|T|}{4n}+O\Big{(}\frac{1}{n}\Big{)}.$
Solving this inequality, we obtain the desired result. ∎
Lemma 3.8.
For each $v\in V(G)$ with $d(v)=\Omega(n)$, $x_{v}=\Theta(n^{-1/2})$.
Proof.
Let $v$ be a vertex with $d(v)=\Omega(n)$. Using the eigenvalue – eigenvector equation with respect to $v$, we conclude that
$(\lambda_{1}^{(\alpha)}-\alpha d(v))x_{v}\geq(1-\alpha)\sum_{u\in N(v)\setminus T}x_{u}\geq(1-\alpha)\cdot\frac{d(v)-O(1)}{2\sqrt{n}}=\Omega(\sqrt{n}).$
The second inequality follows from Lemma 3.7. On the other hand, $\lambda_{1}^{(\alpha)}<n$. Therefore, $x_{v}=\Omega(n^{-1/2})$. Combining with Lemma 3.6, we get the desired result. ∎
Lemma 3.9.
For each $v\in V(G)$ with $d(v)=\Omega(n)$, $|z_{v}|=o(n^{-1/2})$.
Proof.
Suppose on the contrary that there exists a vertex $v$ with $d(v)=\Omega(n)$ such that $|z_{v}|>c/\sqrt{n}$ for some $c>0$. By the eigenvalue – eigenvector equation for $\lambda_{n}^{(\gamma)}$ with respect to $v$, we have
$\bigg{|}\sum_{u\in N(v)}z_{u}\bigg{|}=\frac{\big{|}\big{(}\lambda_{n}^{(\gamma)}-\gamma d(v)\big{)}z_{v}\big{|}}{1-\gamma}=\Omega(\sqrt{n}).$
On the other hand, we see
$\bigg{|}\sum_{u\in N(v)}z_{u}\bigg{|}\leq\sum_{u\in N(v)\cap L}|z_{u}|+\sum_{u\in N(v)\setminus L}|z_{u}|\leq\sqrt{|L|}+\varepsilon\sqrt{n},$
which yields that $|L|=\Omega(n)$ for sufficiently small $\varepsilon$, a contradiction completing the proof. ∎
We are now prepared to show that $G$ contains exactly one vertex of small degree.
Lemma 3.10.
The graph $G$ has exactly one vertex with degree $o(n)$.
Proof.
We prove this lemma by contradiction. Let $M$ be the set of vertices whose degrees are $o(n)$. Then
$2e(G)<|M|\cdot o(n)+(n-|M|)n=n^{2}-|M|(1-o(1))n.$
Moreover, by Lemma 3.2, we have $|M|=O(1)$. Assume that $|M|\geq 2$. In the following we shall get contradiction by considering the cases $|M|\geq 3$ and $|M|=2$, respectively. Together with Lemma 3.5 this implies that $|M|=1$.
To finish the proof, set $p(u,v):=\alpha x_{u}^{2}+2(1-\alpha)x_{u}x_{v}+\alpha x_{v}^{2}$. By Lemma 3.6, $x_{v}<(1+o(1))n^{-1/2}$ for each $v\in V(G)$. Together with (2.2) and Proposition 2.1, this gives
$\lambda_{1}^{(\alpha)}$ $=\sum_{uv\in E(G\setminus M)}p(u,v)+\sum_{uv\in E(M,V(G)\setminus M)\,\cup\,E(M)}p(u,v)$
$<\lambda_{1}^{(\alpha)}(K_{n-|M|})+2(1+o(1))\bigg{(}\sum_{u\in M}d(u)\bigg{)}\cdot\frac{1}{n}$
$<n-|M|-1+o(1),$
where the last equality is due to $|M|=O(1)$. If $|M|\geq 3$, then $\lambda_{1}^{(\alpha)}<n-4+o(1)$, a contradiction to Lemma 3.1.
Now, assume that $|M|=2$. Let $M=\{u,v\}$, and $|z_{u}|\geq|z_{v}|$. We shall show that $\lambda_{n}^{(\gamma)}$ is not excessively small, specifically, $\lambda_{n}^{(\gamma)}>\gamma^{2}(1-\gamma)/12$. Equation (2.3) implies that
$\begin{split}\lambda_{n}^{(\gamma)}&=(2\gamma-1)\sum_{w\in V(G)}d(w)z_{w}^{2}+(1-\gamma)\sum_{ij\in E(G)}(z_{i}+z_{j})^{2}\\ &\geq(1-\gamma)\sum_{ij\in E(G)}(z_{i}+z_{j})^{2}.\end{split}$ (3.1)
On the other hand, in view of Lemma 3.9 we find
$z_{u}^{2}+z_{v}^{2}=1-\sum_{s\in V(G)\setminus\{u,v\}}z_{s}^{2}=1-o(1),$
which yields that
$|z_{u}|>\frac{1}{\sqrt{2}}-o(1).$ (3.2)
We consider the following two cases.
Case 1. $d(u)\geq 2$. Let $w$ be a neighbour of $u$ other than $v$. By Lemma 3.9, $|z_{w}|=o(n^{-1/2})$. Recall that $\frac{1}{2}\leq\gamma<1$. Hence, (3.1) and (3.2) together imply
$\lambda_{n}^{(\gamma)}\geq(1-\gamma)(z_{u}+z_{w})^{2}>\frac{\gamma^{2}(1-\gamma)}{12}.$
Case 2. $d(u)=1$. It is enough to consider the case that $u$ and $v$ are adjacent, otherwise we have $\lambda_{n}^{(\gamma)}>\gamma^{2}(1-\gamma)/12$ using similar arguments as Case 1. Now, let $uv\in E(G)$. By the eigenvalue – eigenvector equation for $\lambda_{n}^{(\gamma)}$ with respect to $u$,
$(\lambda_{n}^{(\gamma)}-\gamma)z_{u}=(1-\gamma)z_{v}.$ (3.3)
Obviously, we may assume that $|z_{v}|>c$ for some small constant $c>0$, otherwise, we have $\lambda_{n}^{(\gamma)}>\gamma-o(1)$ by (3.3). Let $w$ be a neighbor of $v$ other than $u$. It follows from Lemma 3.9 that
$\lambda_{n}^{(\gamma)}\geq(1-\gamma)(z_{v}+z_{w})^{2}>\frac{1-\gamma}{4}z_{v}^{2}.$ (3.4)
From (3.3), we know $z_{v}=\frac{(\lambda_{n}^{(\gamma)}-\gamma)z_{u}}{1-\gamma}$. Combining this with (3.4) yields
$\lambda_{n}^{(\gamma)}>\frac{1-\gamma}{4}z_{v}^{2}>\frac{(\lambda_{n}^{(\gamma)}-\gamma)^{2}}{4(1-\gamma)}z_{u}^{2}>\frac{(\lambda_{n}^{(\gamma)}-\gamma)^{2}}{12(1-\gamma)},$
where the last inequality follows from (3.2). The above inequality is equivalent to
$(\lambda_{n}^{(\gamma)})^{2}-(12-10\gamma)\lambda_{n}^{(\gamma)}+\gamma^{2}<0.$
Solving this inequality, we obtain
$\lambda_{n}^{(\gamma)}$ $>\frac{\gamma^{2}}{6-5\gamma+\sqrt{(6-5\gamma)^{2}-\gamma^{2}}}$
$>\frac{\gamma^{2}}{2(6-5\gamma)}$
$>\frac{\gamma^{2}(1-\gamma)}{12}.$
To sum up, we have $\lambda_{n}^{(\gamma)}>\gamma^{2}(1-\gamma)/12$.
Finally, combining with Lemma 3.1, we find that
$\lambda_{1}^{(\alpha)}$ $=\big{(}\lambda_{1}^{(\alpha)}-\beta\cdot\lambda_{n}^{(\gamma)}\big{)}+\beta\cdot\lambda_{n}^{(\gamma)}$
$>n-2-\beta\gamma+\frac{\beta\gamma^{2}(1-\gamma)}{12}$
$\geq n-3+\frac{\beta\gamma^{2}(1-\gamma)}{12}.$
This is a contradiction with $\lambda_{1}^{(\alpha)}<n-|M|-1+o(1)=n-3+o(1)$. The proof of the lemma is completed. ∎
Hereafter, we assume that $w$ is the unique vertex such that $d(w)=o(n)$.
3.3 Proof of Theorem 1.1
We now combine the results from the previous subsection to prove Theorem 1.1.
Proof of Theorem 1.1..
For any $u,v\in V(G)\setminus\{w\}$, it follows from Lemma 3.8 and Lemma 3.9 that
$\alpha(x_{u}-x_{v})^{2}-\beta\gamma(z_{u}-z_{v})^{2}+2(x_{u}x_{v}-\beta z_{u}z_{v})>0.$
Combining with Lemma 3.3, we deduce that the induced subgraph $G[V(G)\setminus\{w\}]$ is a complete graph. To finish the proof, we show that $d(w)=1$. Indeed, by Lemma 3.9 again,
$z_{w}^{2}=1-\sum_{u\in V(G)\setminus\{w\}}z_{u}^{2}=1-o(1).$
If $d(w)\geq 2$, then deleting one edge between $w$ and $V(G)\setminus\{w\}$ will increase $\lambda_{1}^{(\alpha)}(G)-\beta\cdot\lambda_{n}^{(\gamma)}(G)$ by Lemma 3.3, a contradiction completing the proof. ∎